# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ 3.2 Lagrangian Relaxation with Estimated Constraints
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{T}(\boldsymbol{\pi}-\boldsymbol{\pi}^{{}^{\prime}})=r}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\,.$ (9)
This reduces the inner minimisation problem to a less intensive discrete optimisation, where we only have to search over the neighbouring vertices of the optimal policy in $\hat{\mathcal{F}}$ for a solution. Now, a natural question arises around this formulation:
Can we track $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})$ in the projection lemma over time as we sequentially estimate the constraints?
@@
For a sequence {^$\{\hat{\mathcal{F}}_{t}\}_{t\in\mathbb{N}}$ and $\{\hat{{\boldsymbol{\lambda}}}_{t}\}_{t\in\mathbb{N}}$, we first show that (a) $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$, (b) ${\boldsymbol{\lambda}}^{\star}$ is unique, and (c) $\lim_{t\rightarrow\infty}\hat{{\boldsymbol{\lambda}}}_{t}\rightarrow{\boldsymbol{\lambda}}^{\star}$.
Thus, for any $\boldsymbol{\omega}\in\mathcal{F}$ and ${\boldsymbol{\mu}}$,
where ${\boldsymbol{\lambda}}^{*}$ is such that for any ${\boldsymbol{\lambda}}^{*}\in\operatorname*{arg\,min}_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)$. (Proof is in Appendix D.)
Outer Optimisation Problem. As we guarantee the convergence of $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}}_{t})$ as ^$\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$, we are left with the outer optimisation problem in Equation (1). Since it is a linear problem in $\boldsymbol{\omega}$, we can use a linear programming method which would lead to one of the vertices of ^$\hat{\mathcal{F}}_{t}$. But to be sure of an existence of a solution at each $t\in\mathbb{N}$, we try and derive some well-behavedness properties of the optimal allocation $\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$. First, we observe that our estimates of the mean vector converge to ${\boldsymbol{\mu}}$ as $t\rightarrow\infty$. Hence, we also get $\lim_{t\rightarrow\infty}\mathcal{D}(\boldsymbol{\omega},\hat{{\boldsymbol{\mu}}}_{t},\hat{\mathcal{F}}_{t})\rightarrow\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})$. Now, we ensure well-behavedness and existence of an optimal allocation for all $t>0$.
@@
(Existence of unique optimal allocation) For all ${\boldsymbol{\mu}}\in\mathcal{D}$, $\boldsymbol{\omega}^{\star}({\boldsymbol{\mu}})$ satisfies the conditions
Proof is in Appendix E.
Characterising the Lower Bound for Gaussians. Since we can derive explicit form of the optimisation problem for Gaussian reward distributions, we characterise it further to relate our lower bound with the lower bound for known constraints.
Theorem 4.
Let $\{P_{a}\}_{a\in[K]}$ be Gaussian distributions with equal variance $\sigma^{2}>0$ , $T_{\hat{\mathcal{F}},r}^{-1}$ is
${\max_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\Bigg{\{}\frac{(r-{\boldsymbol{\mu}}^{T}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{2\sigma^{2}\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{T}\tilde{\mathbf{A}}\boldsymbol{\omega}\Bigg{\}}\,,}$
where $\mathrm{Diag}(1/\boldsymbol{\omega}_{a})$ is a $K$-dimensional diagonal matrix with $a$-th diagonal entry $1/\boldsymbol{\omega}_{a}$ and $\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})$ is the set of neighbouring policies of $\boldsymbol{\pi}$ in $\hat{\mathcal{F}}$. (Proof is in Appendix F.)
Now, to connect with the lower bound under known constraints, we fix $r=0$, i.e we search for the ‘true’ optimal policy $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$ rather than an $r$-good feasible policy.
@@
Part i. Then, we get $\frac{2\sigma^{2}K}{C_{\mathrm{known}}}(1+\mathfrak{s}_{\tilde{\mathbf{A}}})\leq T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\leq\frac{2\sigma^{2}K}{C_{\mathrm{known}}}$, where $C_{\mathrm{known}}=\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}d_{\boldsymbol{\pi}^{\prime\prime}}^{2}$.
Part ii. Additionally, $T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\geq\frac{H}{\kappa^{2}}(1+\mathfrak{s}_{\tilde{\mathbf{A}}})$, where $H$ is inversely proportional to the sum of squares of gaps and $\kappa_{\mathrm{known}}$ is the condition number of a sub-matrix of $\mathbf{A}$ consisting ${K}$ linearly independent active constraints for $\boldsymbol{\pi}^{\star}$. (Proof is in Appendix F.1).
Connection to Existing Lower Bounds. (a) Pure exploration under known constraints. The upper and lower bounds on characteristic time coincides with the existing lower bound under known constraints, i.e. when $\epsilon=0$, i.e. $\hat{\mathcal{F}}=\mathcal{F}$. (b) BAI without constraints. In BAI, we consider only deterministic policies (or pure strategies) of playing a single arm. Then, we get $d_{\boldsymbol{\pi}_{a}}=\frac{{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi})_{a}}{|(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi})_{a}|}={\boldsymbol{\mu}}^{\star}-{\boldsymbol{\mu}}_{a}$, i.e. the sub-optimality gap for arm $a$. Here, ${\boldsymbol{\mu}}^{\star}$ is the mean of the best arm. In our setting, if there are no constraints involved then $\tilde{\mathbf{A}}=\mathbf{A}$ and so $\epsilon=0$. Then the lower bound expression in Part ii. of Corollary 1 is inversely proportional to the sum-squared sub-optimal gaps which resonates with the complexity measure in standard BAI with only simplex constraints (Kaufmann et al., (2016)). Though the lower bound stated in Corollary 1 successfully encompasses the effect of unknown linear constraint by introducing novel constraint dependent factors that scales the lower bound.
4 LATS AND LAGEX: ALGORITHM DESIGN AND ANALYSIS
@@
We require two distributional assumptions on rewards and constraints. (i) Reward distributions $\{P_{a}\}_{a=1}^{K}$ are sub-Gaussian one parameter exponential family with mean vector ${\boldsymbol{\mu}}\in\mathcal{D}$. (ii) Each constraint follows a sub-Gaussian ${K}$-parameter exponential family parameterised by $\mathbf{A}^{i}$ for $i\in[d]$.
These assumptions are standard in bandits under constraints (Carlsson et al.,, 2024; Degenne and Koolen,, 2019; Pacchiano et al.,, 2020, 2024).
Algorithm Design. Any algorithm in pure exploration setting comprises of three main components. The first one is a sequential hypothesis testing that decides whether we need to keep sampling or not or popularly known as the "stopping criterion". To come up with such a criterion we need to make sure that we have gathered sufficient information about all the parameters in estimation, specifically ${\boldsymbol{\mu}}$ and $\mathbf{A}$ in our case. So it seems natural to enforce criterion on both mean and constraint concentration.
@@
Thus, while implementing Algorithm 1 and 2, we just need to check the first condition to stop sampling. Refer Appendix G.1 for Proof.
Component 1: Stopping Rule. Stopping rule of an explorer decides when to stop sampling. While exploring, once we gather enough statistical information about the parameters in the system, the test statistic crosses the stopping threshold with the chosen confidence $\delta$, and we stop to recommend the optimal policy.
@@
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum\nolimits_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},\lambda_{a})>\beta(t,\delta)$
where $\beta(t,\delta)\mathrel{\triangleq}3S_{0}\log(1+\log N_{a,t})+S_{0}\mathcal{T}\left(\frac{({K}\wedge d)+\log\frac{1}{\delta}}{S_{0}}\right)$.
Proof in Appendix G.2
@@
1:Input : Time horizon T, $r$, Confidence level $\delta>0,v$
2:Initialization : $\hat{\mathbf{A}}_{0}=\textbf{0}_{{d}\times K}$, $\hat{{\boldsymbol{\mu}}}_{0}=\textbf{0}_{K}$, $\Sigma_{0}=v\mathbf{I}_{K}$, $l_{0}$
3:Play each arm once to set $\mu_{1}$ and $\hat{\mathbf{A}}_{1}$.
4:while $\beta(t-1,\delta)>\mathcal{D}(\boldsymbol{\omega}_{t-1}^{\star},\hat{{\boldsymbol{\mu}}}_{t-1},\hat{\mathcal{F}}_{t-1},\boldsymbol{l}_{t-1}^{\star})$ do $\rightsquigarrow$ Proposition 1
5:     $\boldsymbol{\pi}_{t}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\hat{{\boldsymbol{\mu}}}_{t-1}^{T}\boldsymbol{\pi}$
@@
9:     Feedback : Observe reward $r_{t}$ and cost $\mathbf{A}a_{t}$, and update $\Sigma_{t}$, $\hat{{\boldsymbol{\mu}}}_{t}$ and $\hat{\mathbf{A}}_{t}$
10:     t = t + 1
11:end while
12:Recommended policy:$\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\hat{{\boldsymbol{\mu}}}_{t}^{T}\boldsymbol{\pi}$
Algorithm 2 LAGEX - LAgrangian Game EXplorer
1:Input Time horizon T, $r$, Confidence level $\delta>0,v$
2:Play each arm once to set $\mu_{1}$ and $\hat{\mathbf{A}}_{1}$.
3:while $\beta(t-1,\delta)>\mathcal{D}(\boldsymbol{\omega}_{t-1}^{\star},\hat{{\boldsymbol{\mu}}}_{t-1},\hat{\mathcal{F}}_{t-1},\boldsymbol{l}_{t-1}^{\star})$ do $\rightsquigarrow$ Proposition 1)
@@
5:     Optimal allocation $\boldsymbol{\omega}_{t}^{\star}$rightsquigarrow Using AdaGrad via Theorem 4
@@
8:     Feedback : Observe reward $r_{t}$ and cost $\mathbf{A}a_{t}$, and update $\Sigma_{t}$, $\hat{{\boldsymbol{\mu}}}_{t}$ and $\hat{\mathbf{A}}_{t}$
@@
9:     Compute confusing instance : ${\boldsymbol{\lambda}}_{t}$rightsquigarrow Via Proposition 1 plugging in $\boldsymbol{\omega}_{t}^{\star}$, $\boldsymbol{l}_{t}^{\star}$
@@
10:     Confidence intervals: for all $a\in[K]$ $[\alpha_{t,a},\beta_{t,a}]:\{\zeta:N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{t,a},\zeta)\leq g(t)\}$$U_{t}^{a}=\max\big{\{}\frac{g(t)}{N_{a,t}},d(\alpha_{t,a},{\boldsymbol{\lambda}}_{t,a}),d(\beta_{t,a},{\boldsymbol{\lambda}}_{t,a})\big{\}}$

### v2

@@ 3.2 Relaxation with Estimated Constraints
For any $\boldsymbol{\omega}\in\hat{\mathcal{F}}$ and ${\boldsymbol{\mu}}\in\mathcal{D}$, the Lagrangian relaxation in Equation (6) satisfies $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})=\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{{}^{\prime}})=r}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\,.$
Proposition 1 reduces the inner minimisation problem to a less intensive discrete optimisation, where we only have to search over the neighbouring vertices of the optimal policy in $\hat{\mathcal{F}}$ for a solution. Now, a natural question arises around this formulation:
[…]
To answer this, we first prove convergence of estimated feasible set and alternating instance in order to prove convergence of $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}}_{t})$ in the theorem below. Detailed proof is in Appendix D.
@@
For a sequence $\{\hat{\mathcal{F}}_{t}\}_{t\in\mathbb{N}}$ and $\{\hat{{\boldsymbol{\lambda}}}_{t}\}_{t\in\mathbb{N}}$, we show that (a) $\lim_{t\rightarrow\infty}\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$, (b) $\boldsymbol{\lambda}$$^{\star}$ is unique, and (c) $\lim_{t\rightarrow\infty}\hat{{\boldsymbol{\lambda}}}_{t}\rightarrow{\boldsymbol{\lambda}}^{\star}$. Then, for any $\boldsymbol{\omega}\in\mathcal{F}$ and $\boldsymbol{\mu}$}, $\lim_{t\rightarrow\infty}\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}}_{t})\rightarrow\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F}),$ where ${\boldsymbol{\lambda}}^{*}\in\operatorname*{arg\,min}_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)$.
II. The Outer Optimisation Problem. As we guarantee the convergence of $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}}_{t})$ as $\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$, we are left with the outer optimisation problem in Equation (7). Since it is a linear problem in $\boldsymbol{\omega}$, we can use a linear programming method leading to a vertex of $\hat{\mathcal{F}}_{t}$. But to be sure of an existence of a solution at each $t\in\mathbb{N}$, we need well-behavedness properties of the optimal allocation $\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$. First, we observe that our estimates of the mean vector converge to $\boldsymbol{\mu}$} as $t\rightarrow\infty$. Hence, we also get $\lim_{t\rightarrow\infty}\mathcal{D}(\boldsymbol{\omega},\hat{{\boldsymbol{\mu}}}_{t},\hat{\mathcal{F}}_{t})\rightarrow\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})$. Now, we ensure well-behavedness and existence of an optimal allocation for all $t>0$.
Theorem 3 (Existence of unique optimal allocation).
For all ${\boldsymbol{\mu}}\in\mathcal{D}$, $\boldsymbol{\omega}^{\star}({\boldsymbol{\mu}})$ satisfies: 1. Both the sets $\hat{\mathcal{F}}$ and $\boldsymbol{\omega}^{\star}({\boldsymbol{\mu}})$ are closed and convex. 2. For all ${\boldsymbol{\mu}}\in\mathcal{D}$ and $\boldsymbol{\omega}\in\hat{\mathcal{F}}$, $\lim_{t\rightarrow\infty}\mathcal{D}(\boldsymbol{\omega},\hat{{\boldsymbol{\mu}}}_{t},\hat{\mathcal{F}}_{t})$ is continuous. 3. Reciprocal of the characteristic time $\lim_{t\rightarrow\infty}T^{-1}_{\hat{\mathcal{F}}_{t},r}({\boldsymbol{\mu}})$ is continuous for all ${\boldsymbol{\mu}}\in\mathcal{D}$. 4. For all ${\boldsymbol{\mu}}\in\mathcal{D}$, ${\boldsymbol{\mu}}\rightarrow\boldsymbol{\omega}^{\star}({\boldsymbol{\mu}})$ is upper hemi-continuous. Thus, the optimization problem $\max_{\boldsymbol{\pi}\in\hat{{\mathcal{F}}}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$ has a unique solution.
Characterising Lower Bound for Gaussians. Since we can derive explicit form of the optimisation problem for Gaussian reward distributions, we characterise it further to relate our lower bound with the lower bound for known constraints (Appendix F).
@@
Let $\{P_{a}\}_{a\in[K]}$ be Gaussian distributions with equal variance $\sigma^{2}>0$ , $T_{\hat{\mathcal{F}},r}^{-1}$ is $\max_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\min_{\boldsymbol{\pi}^{\prime}\in\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})}$ $\Bigg\{\frac{(r-{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}))^{2}}{2\sigma^{2}\|\boldsymbol{\pi}-\boldsymbol{\pi}^{\prime}\|_{\mathrm{Diag}(1/\boldsymbol{\omega}_{a})}^{2}}-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\Bigg\}\,$, where $\mathrm{Diag}(1/\boldsymbol{\omega}_{a})$ is a $K$-dimensional diagonal matrix with $a$-th diagonal entry $1/\boldsymbol{\omega}_{a}$ and $\nu_{\hat{\mathcal{F}}}(\boldsymbol{\pi})$ is the set of neighbouring policies of $\boldsymbol{\pi}$ in $\hat{\mathcal{F}}$.
Now, to connect with the lower bound under known constraints, we fix $r=0$, i.e we search for the ‘true’ optimal policy $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$ rather than an $r$-optimal feasible policy.
@@
Part i. Then, we get
$\frac{2\sigma^{2}K}{C_{\mathrm{known}}}(1+\mathfrak{s}_{\tilde{A}})\leq T_{\hat{\mathcal{F}},0}({\boldsymbol{\mu}})\leq\frac{2\sigma^{2}K}{C_{\mathrm{known}}},$
where $C_{\mathrm{known}}=\min_{\boldsymbol{\pi}^{\prime\prime}\in\nu_{\mathcal{F}}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})}d_{\boldsymbol{\pi}^{\prime\prime}}^{2}$.
Part ii.
[…]
where $\mathfrak{s}_{\tilde{A}}\mathrel{\triangleq}\frac{\max_{i\in[1,d]}\tilde{\mathbf{A}}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}}{\min_{i\in[1,d]}\tilde{\mathbf{A}}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}}$ is the estimated shadow price and $H$ is inversely proportional to the sum of squares of gaps and $\kappa_{\mathrm{known}}$ is condition number of a sub-matrix of $A$ with K{K} linearly independent active constraints for $\boldsymbol{\pi}^{\star}$.
Remark 3 (Connection to Existing Lower Bounds.).
(a) Pure exploration under known constraints. The upper and lower bounds on characteristic time coincides with the existing lower bound under known constraints, i.e. when $\epsilon=0$, i.e. $\hat{\mathcal{F}}$$\mathcal{F}$. (b) BAI without constraints. In BAI, we consider only deterministic policies (or pure strategies) of playing a single arm. Then, we get $d_{\boldsymbol{\pi}_{a}}=\frac{{\boldsymbol{\mu}}^{\top}(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi})_{a}}{|(\boldsymbol{\pi}^{\star}_{\mathcal{F}}-\boldsymbol{\pi})_{a}|}={\boldsymbol{\mu}}^{\star}-{\boldsymbol{\mu}}_{a}$, i.e. the sub-optimality gap for arm $a$. Here, $\boldsymbol{\mu}$$^{\star}$ is the mean of the best arm. In our setting, if there are no constraints involved then $\tilde{A}=A$ and so $\epsilon=0$. Then the lower bound expression in Part ii. of Corollary 1 is inversely proportional to the sum-squared sub-optimal gaps which resonates with the complexity measure in standard BAI with only simplex constraints (Kaufmann et al., (2016)). Though the lower bound stated in Corollary 1 successfully encompasses the effect of unknown linear constraint by introducing novel constraint dependent factors that scales the lower bound.
4 Algorithm Design
@@
We require two distributional assumptions on rewards and constraints. (i) Reward distributions $\{P_{a}\}_{a=1}^{K}$ are sub-Gaussian one parameter exponential family with mean vector ${\boldsymbol{\mu}}\in\mathcal{D}$. (ii) Each constraint follows a sub-Gaussian K{K}-parameter exponential family parameterised by $A^{i}$ for $i\in[d]$.
These distributional assumptions are standard in bandits under constraints (Carlsson et al.,, 2024; Degenne and Koolen,, 2019; Pacchiano et al.,, 2020, 2024).
Algorithm Design. Any algorithm in pure exploration setting consists of three main components. The first one is a sequential hypothesis testing that decides whether we need to keep sampling or not, popularly known as the “stopping criterion”. To come up with such a criterion we need to make sure that we have gathered sufficient information about all the parameters in estimation, specifically $\boldsymbol{\mu}$} and $A$ in our case. So it seems natural to enforce criterion on both mean and constraint concentration.
@@
Thus, while implementing Algorithm 2 and 3, we just need to check the first condition to stop sampling.
Component 1: Stopping Rule. During exploration, once we gather enough statistical information about the parameters in the system, the test statistic crosses the stopping threshold with the chosen confidence $\delta$, and we stop to recommend the optimal policy.
@@
The Chernoff stopping rule to ensure $(1-\delta)$-correctness and $(1-\delta)$-feasibility is $\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum\nolimits_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\beta(t,\delta)\,$, where $\beta(t,\delta)\mathrel{\triangleq}3S_{0}\log(1+\log N_{a,t})+S_{0}\mathcal{T}\left(\frac{({K}\wedge d)+\log\frac{1}{\delta}}{S_{0}}\right)$ and $0\leq S_{0}\leq{K}$.
@@
1:Input : Tolerance $r>0$, Confidence level $\delta>0$
2:Initialization : $\hat{A}_{0}=\textbf{0}_{{d}\times K}$, $\hat{{\boldsymbol{\mu}}}_{0}=\textbf{0}_{K}$, $\Sigma_{0}=v\mathbf{I}_{K}$, $l_{0}$
3: Play each arm once to set ${\boldsymbol{\mu}}_{1}$ and $\hat{A}_{1}$.
4:while $\beta(t-1,\delta)>\mathcal{D}(\boldsymbol{\omega}_{t-1}^{\star},\hat{{\boldsymbol{\mu}}}_{t-1},\hat{\mathcal{F}}_{t-1},\boldsymbol{l}_{t-1}^{\star})$ do
5:  $\boldsymbol{\pi}_{t}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\hat{{\boldsymbol{\mu}}}_{t-1}^{\top}\boldsymbol{\pi}$
@@
9:  Feedback : Observe reward $r_{t}$ and cost $\mathbf{c}_{t}$, and update $\Sigma_{t}$, $\hat{{\boldsymbol{\mu}}}_{t}$ and $\hat{A}_{t}$
10:end while
11:Recommended policy: $\hat{{\boldsymbol{\pi}}}=\operatorname*{arg\,max}\limits_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}$
Component 2: Recommendation rule. Once the stopping rule is fired, the agent recommends a policy based on the current estimate of $\hat{{\boldsymbol{\mu}}}_{t}$ according the rule $\hat{{\boldsymbol{\pi}}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{\tau_{\delta}}}^{r}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}.$
LATS.
The algorithm LATS (Algorithm 2) uses the Track and Stop strategy adapted to the unknown constraint setting. The algorithm warms up the parameter estimates by playing each arm once. Until the stopping rule is fired, it computes the optimal allocation (Line 4) under the estimated feasible space by solving the Lagrangian relaxed optimization problem in Proposition 1. By plugging the optimal allocation we perform a bounded optimisation on the Lagrangian multiplier using Theorem 1. The algorithm further uses C-tracking (Garivier and Kaufmann,, 2016) to track actions taken per step. Finally, we observe the instantaneous reward and cost feedback to update the parameter estimates (Line 9).
Theorem 6.
Let $\mathfrak{s}$ be the shadow price $\mathfrak{s}\mathrel{\triangleq}\frac{\Gamma_{\mathrm{max}}}{\Gamma_{\mathrm{min}}}=\frac{\max_{i\in[1,d]}(-A\boldsymbol{\pi}^{\star}_{\mathcal{F}})}{\min_{i\in[1,d]}(-A\boldsymbol{\pi}^{\star}_{\mathcal{F}})}$ of the slack $\Gamma$. Under Assumption 1 and 2, the expected stopping time of LATS satisfies $\lim_{\begin{subarray}{c}\delta\rightarrow 0\end{subarray}}\frac{\mathbb{E}[\tau_{\delta}]}{\log(1/\delta)}\leq\alpha T_{\mathcal{F},r}({\boldsymbol{\mu}})(1+\mathfrak{s})$ for any $\alpha>1$.
Implications. Theorem 6 suggests that LATS (Algorithm 2) is asymptotically optimal up to problem dependent constants. This upper bound captures the effect of unknown linear constraints in shadow price, novel in the constrained bandit literature. If any one of the constraints in the bandit environment becomes more and more sensitive, minimum slack decreases, consequently $\mathfrak{s}$ increases and the identification problem becomes harder to solve. Thus, $\mathfrak{s}$ behaves as a sensitivity or stability parameter that explains the hardness of the pure exploration problem in hand through the structure of the true constraints.
Algorithm 3 LAGEX- LAgrangian Gamified EXplorer
1:Input : Tolerance $r>0$, Confidence level $\delta>0$
2: Play each arm once to set ${\boldsymbol{\mu}}_{1}$ and $\hat{A}_{1}$.
3:while $\beta(t-1,\delta)>\mathcal{D}(\boldsymbol{\omega}_{t-1}^{\star},\hat{{\boldsymbol{\mu}}}_{t-1},\hat{\mathcal{F}}_{t-1},\boldsymbol{l}_{t-1}^{\star})$ do
@@
5:  Optimal allocation $\boldsymbol{\omega}_{t}^{\star}$ $\rightsquigarrow$ Using AdaGrad via Theorem 4
@@
8:  Feedback : Observe reward $r_{t}$ and cost $\mathbf{c}_{t}$, and update $\Sigma_{t}$, $\hat{{\boldsymbol{\mu}}}_{t}$ and $\hat{A}_{t}$
@@
9:  Compute confusing instance : ${\boldsymbol{\lambda}}_{t}$ $\rightsquigarrow$ Via Proposition 1 plugging in $\boldsymbol{\omega}_{t}^{\star}$, $\boldsymbol{l}_{t}^{\star}$
@@
10:  Confidence intervals: for all $a\in[K]$ $\alpha_{t,a},\beta_{t,a}]:\{\zeta:N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{t,a},\zeta)\leq g(t)\}$$U_{t}^{a}=\max\big\{\frac{g(t)}{N_{a,t}},d(\alpha_{t,a},{\boldsymbol{\lambda}}_{t,a}),d(\beta_{t,a},{\boldsymbol{\lambda}}_{t,a})\big\}$